# The initial value problem in cosmology: An alternative derivation for beginners

### v1

@@ 4.2 Choosing the pressure equation and the continuity equation
where we have explicit appearance of $\dot{\rho}/\dot{a}$. The above equation always remain well defined as because $\dot{a}$ never vanishes in the time interval $[t_{i}\,,\,t_{f}]$. Multiplying all the terms of the above equation by $\dot{a}/a$ we get
@@
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $[t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero at any point inside the interval. For any initial condition $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$ we can always solve the continuity equation to get a positive $\rho(t)$ for the cases where $\omega=0$ or $\omega=1/3$. Using the expression of $\rho(t)$ in the pressure equation we can get $a(t)$. Although we have produced the dynamics from an arbitrary, but otherwise valid, initial condition, this dynamical evolution will not satisfy the Friedmann equation. The resultant dynamical evolution cannot be a solution to the Einstein equation.
@@ 4.3 Choosing the acceleration equation and the continuity equation
This equation is also well defined throughout the interval $[t_{i}\,,\,t_{f}]$ as $\dot{a}$ never vanishes inside the interval. Multiplying both sides of the above equation by $2\dot{a}/a$ we get
@@
This equation is similar to Eq. (23) but not exactly the same. The last equation is difficult to interpret as we still cannot fix the value of $\tilde{k}$ from our chosen set of equations. In such a case we cannot even guess whether the Friedmann equation is valid. Even we do not know from our set of equations whether the pressure equation is valid. For some arbitrary initial conditions we can generate dynamical evolution from our set of equations but we will not be able to know whether our solution satisfies the Friedmann equation and the pressure equation. The situation is more serious than the case before. If we start with any arbitrary initial conditions we do not even know whether the Einstein equation is valid.
@@ 4.4 The link between all the choices is in the initial condition
From our previous discussion we have seen that all the pairs of chosen differential equations may not reproduce the other differential equations. In such cases one can have a resultant dynamics which is not a unique solution of the Einstein equation. All the previous discussion was based on arbitrary choice of initial conditions. We did not mention any scheme for comparing initial conditions. As because our system admits both first order and second order differential equations, it is difficult to give an ansatz for equivalence of initial condition. Equivalence of initial condition means an initial condition which produces a proper solution to the Einstein equation, no matter whichever pair of equations one works with. When we use such an initial condition, we demand all the chosen pairs of differential equations will produce the same time development of the system and all the differential equations will be simultaneously satisfied at all instants of time inside the chosen interval $[t_{i}\,,\,t_{f}]$. The pressure equation and continuity equation as well as the acceleration equation and the continuity equation requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. There is no unique way to choose this triplet of numbers. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. Apparently it seems it is impossible to specify the initial condition which can be applied to all the three pairs of equations. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
@@
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. From our previous discussion we have seen that when we produce the dynamics from the pressure equation and the continuity equation from any arbitrary initial condition the Friedmann equation may not hold. But once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $[t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (23). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As a consequence $F(t)$ is always zero inside the interval $[t_{i}\,,\,t_{f}]$. This is a particularly interesting phenomena. The initial constraint condition which produces the initial values of the variables, becomes a dynamical entity and, remains dynamically true throughout the relevant time interval. The constraint propagates through time. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval. The generated dynamics is naturally a solution of the Friedmann equation. Once the initial condition is chosen from the Friedmann constraint, the resulting dynamical solution satisfies all the relevant dynamical equations, simultaneously, in $[t_{i}\,,\,t_{f}]$.
@@
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. In this case we do not know the value of $\tilde{k}$. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then what happens? We see from Eq. (29), this initial condition can only be chosen when $\tilde{k}=k$. In this case, the Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (22). In this case also we see, from Eq. (29), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $[t_{i}\,,\,t_{f}]$. It is interesting to note that although the chosen pair of dynamical equations did not have any information about $k$, appearing in the Friedmann equation, the initial condition passes on the information and this information is not lost as the initial constraint propagates through time. The result shows that if we choose the initial condition through the Friedmann constraint then the Friedmann equation always remains true. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
@@
From the previous discussion it is seen that the Friedmann equation acts as a constraint equation, based on which the initial condition of cosmological dynamics is obtained. This constraint equation remains valid throughout the interval $[t_{i}\,,\,t_{f}]$ and consequently it can also be interpreted as a dynamical equation. The dual roles do not contradict because one can choose any time instant to be the initial time $t_{i}$. As a consequence the Friedmann equation becomes a dynamic constraint.
5 The cases where $\dot{a}(t_{0})=0$ and $t_{0}$ is inside the interval $[t_{i}\,,\,t_{f}]$
@@
Even when both $\dot{\rho}$ and $\dot{a}$ tends to zero, inside the interval $[t_{i}\,,\,t_{f}]$, their ratio remains finite.
@@ 5.1 When $t_{0}=t_{i}$
In this case we have $\dot{a}(t_{i})=0$ at the initial point. The Friedmann equation and the continuity equation can generate the dynamics from this point. The discussion in subsection 4.1 showed us that the Friedmann equation and the continuity equation naturally implies the pressure equation as long as $\dot{a}(t)\neq 0$ inside $[t_{i}\,,\,t_{f}]$. Although the pressure equation connot be connected to the Friedmann equation and the continuity equation at $t=t_{i}$, one can still have a specific solution to the full Einstein equation. That a solution exists in the present case can be shown in the following way. We know that the Friedmann equation and the continuity equation can generate a continuous $a(t)$ and $\rho(t)$ in the time interval $[t_{i}\,,\,t_{f}]$. Suppose instead of $t_{i}$, we start at $t_{i}+\epsilon$, where $\epsilon$ is an arbitrarily small positive constant with dimension of time. We assume that $\dot{a}(t_{i})=0$ and consequently $\dot{a}$ never vanishes inside the interval $[t_{i}+\epsilon\,,\,t_{f}]$. The initial condition at $t=t_{i}+\epsilon$ are now generated from the known solution coming out from the Friedmann equation and continuity equation, which gave the dynamical solution in the full time interval $[t_{i}\,,\,t_{f}]$. All the generated solutions for different (but arbitrarily small) $\epsilon$ smoothly matches with the original solution defined in the full time interval for regions $t_{i}+\epsilon\leq t\leq t_{f}$. From our discussion in subsection 4.2 we must conclude that inside the interval $[t_{i}+\epsilon\,,\,t_{f}]$ we must have a proper solution to the Einstein equation as the pressure equation and the continuity equation will also have the same solution produced from the Friedmann equation and continuity equation if the initial condition is obtained from the Friedmann constraint. Now one can make $\epsilon$ arbitrarily small and for each small value of $\epsilon$ we always obtain a unique cosmological solution. Continuity of our generated solution implies that in the limit $\epsilon\to 0$ we will also have a unique solution of the Einstein equation. The solutions from the two different pairs coincide at $t_{i}=t_{0}$ as the pressure equation and the continuity equation starts with initial conditions following the Friedmann constraint.
For the other cases related to Eq. (23) and Eq. (31), we see that if $\dot{a}(t_{i})=0$, it seems that both these equations can be satisfied for any $F(t_{i})$ as long as $\dot{F}(t_{i})=0$. But this arbitrariness disappears once we use the Friedmann constraint initially and fix $F(t_{i})=0$. As a result of this the Friedmann equation will be valid for all points inside the relevant time interval and we will always get a definite solution to the Einstein equation.
@@ 5.2 When $t_{i}<t_{0}<t_{f}$
Next we discuss the case where $\dot{a}(t_{0})=0$ where $t_{0}$ is an intermediate point in the relevant time interval. In this case we can obtain a solution to the Einstein equation if our initial condition satisfies the Friedmann constraint. We can obtain the collapsing solution in the interval $[t_{i}\,,\,t_{f}]$ by using the Friedmann equation and the continuity equation. But is this solution also obtained by the other pairs of equations? Before we proceed we observe that all the relevant dynamical equations, namely the Friedmann equation, pressure equation, acceleration equation and the continuity equation are symmetric around $t=0$, by which we mean
@@
[…]

### v2

@@ 3.2 Choosing the pressure equation and the continuity equation
where we have explicit appearance of $\dot{\rho}/\dot{a}$. The above equation always remain well defined as because $\dot{a}$ never vanishes in the time interval $t_{i}\,,\,t_{f}]$. Multiplying all the terms of the above equation by $\dot{a}/a$ we get
@@
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero at any point inside the interval. Consequently for arbitrary initial conditions this dynamical evolution will not satisfy the Friedmann equation. The resultant dynamical evolution cannot be a valid solution of the cosmological evolution equations.
@@ 3.3 Choosing the acceleration equation and the continuity equation
This equation is also well defined throughout the interval $t_{i}\,,\,t_{f}]$ as $\dot{a}$ never vanishes inside the interval. Multiplying both sides of the above equation by $2\dot{a}/a$ we get
@@
This equation is similar to Eq. (14) but not exactly the same. The last equation is difficult to interpret as we still cannot fix the value of $\tilde{k}$ from our chosen set of equations. In such a case we cannot even guess whether the Friedmann equation is valid.
@@ 3.4 The link between all the choices is in the initial condition
From our previous discussion we have seen that all the pairs of chosen differential equations may not reproduce the other differential equations. All the previous discussion was based on arbitrary choice of initial conditions. We did not mention any scheme for comparing initial conditions. The pressure equation and continuity equation as well as the acceleration equation and the continuity equation requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. There is no unique way to choose this triplet of numbers. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. Apparently it seems it is impossible to specify the initial condition which can be applied to all the three pairs of equations. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
@@
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. Once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (14). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval. The generated dynamics is naturally a solution of the Friedmann equation.
@@
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. In this case we do not know the value of $\tilde{k}$. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then what happens? We see from Eq. (20), this initial condition can only be chosen when $\tilde{k}=k$. In this case, the Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (13). In this case also we see, from Eq. (20), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $t_{i}\,,\,t_{f}]$. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
@@
From the previous discussion it is seen that the Friedmann equation acts as a constraint equation, based on which the initial condition of cosmological dynamics is obtained. This constraint equation remains valid throughout the interval $t_{i}\,,\,t_{f}]$ and consequently it can also be interpreted as a dynamical equation. The dual roles do not contradict because one can choose any time instant to be the initial time $t_{i}$. As a consequence the Friedmann equation becomes a dynamic constraint.
4 The cases where $\dot{a}(t_{0})=0$ and $t_{0}$ is inside the interval $t_{i}\,,\,t_{f}]$
@@
Even when both $\dot{\rho}$ and $\dot{a}$ tends to zero, inside the interval $t_{i}\,,\,t_{f}]$, their ratio remains finite.
@@ 4.1 When $t_{0}=t_{i}$
In this case we have $\dot{a}(t_{i})=0$ at the initial point. That a solution exists in the present case can be shown in the following way. We know that the Friedmann equation and the continuity equation can generate a continuous $a(t)$ and $\rho(t)$ in the time interval $t_{i}\,,\,t_{f}]$. Suppose instead of $t_{i}$, we start at $t_{i}+\epsilon$, where $\epsilon$ is an arbitrarily small positive constant with dimension of time. We assume that $\dot{a}(t_{i})=0$ and consequently $\dot{a}$ never vanishes inside the interval $t_{i}+\epsilon\,,\,t_{f}]$. The initial conditions at $t=t_{i}+\epsilon$ are now generated from the known solution coming out from the Friedmann equation and continuity equation. All the generated solutions for different (but arbitrarily small) $\epsilon$ (possibly using other pairs of differential equations) smoothly matches with the solution defined in the full time interval for regions $t_{i}+\epsilon\leq t\leq t_{f}$. One can make $\epsilon$ arbitrarily small and continuity of our generated solution implies that in the limit $\epsilon\to 0$ we will also have a unique solution of the Einstein equation. The solutions from the different pairs coincide at $t_{i}=t_{0}$ as the pressure equation and the continuity equation starts with initial conditions following the Friedmann constraint.
For the other cases related to Eq. (14) and Eq. (22), we see that if $\dot{a}(t_{i})=0$, it seems that both these equations can be satisfied for any $F(t_{i})$ as long as $\dot{F}(t_{i})=0$. But this arbitrariness disappears once we use the Friedmann constraint initially and fix $F(t_{i})=0$. As a result of this the Friedmann equation will be valid for all points inside the relevant time interval (if we start from pairs of equations not containing the Friedmann equation) and we will always get a definite solution to the Einstein equation.
@@ 4.2 When $t_{i}<t_{0}<t_{f}$
Next we discuss the case where $\dot{a}(t_{0})=0$ where $t_{0}$ is an intermediate point in the relevant time interval. In this case we can obtain a solution to the Einstein equation if our initial condition satisfies the Friedmann constraint. We can follow a solution of the Einstein’s equation involving gravitational expansion and a subsequent gravitational contraction in the interval $t_{i}\,,\,t_{f}]$ by using the Friedmann equation and the continuity equation. The turn-around happens at $t=t_{0}$ when $\dot{a}(t_{0})=0$. Is this solution also obtained by the other pairs of equations? Before we proceed we observe that all the relevant dynamical equations, namely the Friedmann equation, pressure equation, acceleration equation and the continuity equation are symmetric around $t=0$, by which we mean
@@
when $\dot{a}(t=0)=0$. We will use these properties and shift the origin of time to $t=t_{0}$ so that what used to be called $t_{0}$ becomes the origin of time $t=0$. The initial time interval now becomes $t_{i}-t_{0}\,,\,t_{f}-t_{0}]$. In shifted time we have $\dot{a}(0)=0$ which implies a symmetric gravitational collapse. The initial conditions can be well specified in the shifted time as $\rho(0)$ and $a(0)$. Now using any pair of differential equations we can generate the solution of the Einstein equation in the expanding phase, in the interval $t_{i}-t_{0},0]$. In this interval all the solutions will agree if all of them satisfy the Friedmann constraint initially. The symmetry of the Einstein equations now guarantee that these solutions must also agree on the contracting phase. If one now wants to go back to the old time interval, one has to just shift the time origin back to $t_{i}$.
5 On the logical consistency of the theory
From our previous discussion it is seen that one can model cosmological dynamics from any relevant pair of equations (presented in this article) and even if those equations do not involve the Friedmann equation specifically, one cannot forget about the Friedmann equation. All cosmological initial conditions must satisfy the Friedmann equation. Although we showed the relevance of this statement in a simple case, the last statement is valid in more general cases as will be shown later in this article.
The results of this article are related to the initial value problem and its solution in cosmology. Can one set up proper initial conditions in GR and solve the Einstein equations consistently in the future, where one is using the FLRW spacetime? This is the initial value problem in cosmology and it is known that the answer is in the affirmative. A formal discussion on the initial value problem in cosmology and its solution can be found in Refs. [12, 13]. In the rest of this section we discuss about the logical consistency of cosmological dynamics in GR.
Till now what has been discussed indicates that the complexity of the situation is primarily related to the availability of both first order and second order differential equations. Naturally the question arises: Is this a peculiarity of the cosmological equations in GR? In this section we will show that the dynamical structure of the cosmological theory, as predicted by GR, is logically consistent and makes perfect sense.
To understand the logical consistency in the present case, we will like to specify here that some form of a first order differential equation is inevitable in the present case if we have to make the theory meaningful. If we leave out quantum effects, we may safely assume that there cannot be matter creation from the gravitational field. In such a case we must have some condition which predicts that the evolution equation of energy density is a first order differential equation and its form must be $\dot{\rho}\propto\rho$. This is a very general requirement, as for such a relation if $\rho$ is zero initially then $\dot{\rho}$ is also zero initially and hence the matter energy density must remain zero for all times. A second order differential equation in $\rho(t)$ will never satisfy the general requirement, as in this case even if initially $\rho$ and $\dot{\rho}$ is zero, still $\rho(t)$ can attain finite non-zero values in the future.
As the Einstein equations (Friedmann and pressure equations) in cosmology do not contain a time derivative of the energy density of the fluid so those equations does not generate naturally a differential equation for $\rho(t)$, but we know from general assumptions, about the system, that such an equation is inevitable. In such a case there is only one way the first order differential equation for $\rho(t)$ can appear, that intended equation must be derivable from the available Einstein equations. In reality it happens like that as because the continuity equation is obtained from the Friedmann and pressure equations. To obtain the continuity equation one must have to differentiate the Friedmann equation and consequently the Friedmann equation must be a first order differential equation. The Friedmann equation cannot be a second order differential equation as in that case a differentiation of this equation will produce a third order derivative of the scale-factor in the resulting equation which cannot be cancelled by any terms in the other Einstein equation for pressure. Hence the equation giving the energy density of the fluid in terms of the functions of the scale-factor and its derivatives must have to be a first-order differential equation of $a(t)$. As because GR, as a well defined field theory, contains second order derivative terms of the scale-factor, the other equation for pressure must have to be a second order differential equation for $a(t)$.
At this point we must point out that with all the manipulations, as specified in the previous paragraphs, one in general always obtains the desired first order differential equation for $\rho(t)$. This is not a coincidence, rather the theory is formulated in such a way that the desired differential equation is produced due to a tensor identity, called the (second) Bianchi identity, constraining the Einstein tensor. The Bianchi identity naturally makes the theory consistent by introducing the energy conservation condition as predicted by the continuity equation.
The previous discussion made it clear that in GR, the Bianchi identity and the Einstein equation both are responsible for the dynamical structure of the cosmological theory and together they affect the initial condition of the problem. Interestingly, whatever we have discussed about cosmological dynamics in GR can be generalized to a modern modified theory of gravity. In the next section we generalize our results in $f(R)$ gravity.
6 On the redundancy of dynamical equations and initial conditions in $f(R)$ cosmology
Till now we have worked in the ambit of GR and Newtonian cosmology. In this section we will show that the techniques we have derived can be generalized in a modified theory of gravity, called the metric $f(R)$ theory. This modification of gravity is a modern development which became useful in theories of inflation and tackling the dark energy problem. The two reviews in Refs. [14, 15] will help the reader to grasp the essence of the theory and we will closely follow the conventions and notations used in those reviews.
In metric $f(R)$ gravity theory the gravitational action changes from
$S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,R+S_{m}\to S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,f(R)+S_{m}\,$,
where $R$ is the Ricci scalar and $S_{m}$ is the action for the non-gravitational degrees of freedom, generally called the matter action. Here $f(R)$ is a general function of $R$ and we do not choose any particular form of $f(R)$ here as our analysis will be true for any general $f(R)$. The field equations corresponding to the Eqs. (3), (4) in $f(R)$ gravity turns out to be:
$\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $\frac{8\pi G}{3f^{\prime}(R)}(\rho+\rho_{\rm{eff}})\,$, (24)
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $-\frac{8\pi G}{f^{\prime}(R)}(P+P_{\rm{eff}})\,$, (25)
where a prime denotes a derivative with respect to the Ricci scalar, $R$, and
$8\pi G\rho_{\rm{eff}}$ $=$ $\frac{Rf^{\prime}(R)-f(R)}{2}-3H\dot{R}f^{\prime\prime}(R)\equiv\tilde{\rho}_{\rm{eff}}\,$, (26)
$8\pi GP_{\rm{eff}}$ $=$ $\dot{R}^{2}f^{\prime\prime\prime}(R)+2H\dot{R}f^{\prime\prime}(R)+\ddot{R}f^{\prime\prime}(R)-\frac{Rf^{\prime}(R)-f(R)}{2}\equiv\tilde{P}_{\rm{eff}}\,$, (27)
where we have used the Hubble parameter, $H\equiv\dot{a}/a$, to compactly write the above two expressions. The Ricci scalar $R=6[(\ddot{a}/{a})+(\dot{a}^{2}/{a^{2}})+({k}/{a^{2}})]=6[\dot{H}+2H^{2}+(k/a^{2})]$ for the FLRW metric given in Eq. (1). Here $\rho_{\rm{eff}},P_{\rm{eff}}$ are effectively acting like the energy density and the pressure induced by the Ricci curvature of the spacetime. In the above equations $\rho,P$ are the normal energy density and pressure of a barotropic matter fluid present in the FLRW spacetime. We name Eq. (24) as the modified Friedmann equation and Eq. (25) as the modified pressure equation. One can also combine the field equations to write down the modified acceleration equation in the present case. Although we name the relevant field equations as we did before, one must remember that these equations are very different from the corresponding equations in GR as here the modified Friedmann equation is a third order differential equation (apparent from the expression for $\rho_{\rm eff}$), the modified pressure equation and the modified acceleration equation are fourth order differential equations (apparent from the expression for $P_{\rm eff}$).
Remarkably, in the present case if one differentiates the modified Friedmann equation and combines the result with the pressure equation, one gets back the continuity equation for the barotropic fluid, as given in Eq. (6). Consequently, in the present case we have four differential equations of three different orders: one is a third order equation, two are fourth order equations and the other one is a first order equation. The disparity in the degrees of the dynamical equations has increased in the present case. Interestingly, with all these drastic differences with GR, we will show that most of the techniques used to understand the redundancy of the equations are still applicable in the present case and the special role of the modified Friedmann equation remains intact in $f(R)$ gravity.
The above equations show that $f^{\prime}(R)$ modifies the gravitational coupling terms from $G\to G/f^{\prime}(R)$, consequently for gravitational stability of the theory we demand $f^{\prime}(R)>0$. Moreover, it was noted that if $f^{\prime\prime}(R)$ is not positive, one can have another kind of instability of the theory, called the Dolgov-Kawasaki instability [16]. As a consequence of these facts, we will always assume
$f^{\prime}(R)>0\,,\,\,\,\,\,\,\,\,f^{\prime\prime}(R)>0\,$, (28)
in $f(R)$ cosmology. In the field equations of $f(R)$ gravity we have set the cosmological constant, $\Lambda=0$, as $f(R)$ theory can dynamically generate dark energy and we may not explicitly require a cosmological constant in this framework of gravity.
In the present case one can differentiate the modified Friedmann equation with respect to time and combine the result with the continuity equation to obtain the pressure equation. We do not present this derivation as it is a straightforward calculation. We show how one obtains the modified Friedmann constraint from the modified pressure equation and the continuity equation. This calculation will make it clear how the previous concepts can be generalized in the present case. If we choose Eq. (25) and the continuity equation as our basic dynamical equations then we can start to manipulate them in a way done previously. Using the continuity equation we write Eq. (25) as:
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}+\frac{8\pi G}{f^{\prime}}P_{\rm eff}=\frac{8\pi G}{f^{\prime}}\left(\frac{1}{3}\dot{\rho}\frac{a}{\dot{a}}+\rho\right)\,.$ (29)
Multiplying all the terms of the above equation by $\dot{a}/a$ we get
$2\frac{\ddot{a}\dot{a}}{a^{2}}+\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$ (30)
The above equation can also be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}\right)+\frac{3\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$
Subtracting
$3\left(\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}\right)$
from both sides of the last equation, we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\frac{d\rho}{dt}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}\rho-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,$, (31)
which can also be written as:
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}$ $=$ $\frac{d}{dt}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\right]-\frac{8\pi G}{3f^{\prime}}\frac{d\rho_{\rm eff}}{dt}+\frac{8\pi G}{3f^{\prime\,2}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}$ (32)
- $\frac{8\pi G}{f^{\prime}}H\rho_{\rm eff}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,.$
Defining
$Q(t)\equiv\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)-\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\,$, (33)
we can write Eq. (32) as:
$\frac{dQ}{dt}+3HQ=-\frac{1}{3f^{\prime}}\left[\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})-3\left\{\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}\right\}\right]\,.$ (34)
In the present case one can show that:
$\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})=\dot{R}f^{\prime\prime}\left(\frac{1}{2}R-3\dot{H}-3H^{2}\right)=3\dot{R}f^{\prime\prime}\left(H^{2}+\frac{k}{a^{2}}\right)\,$, (35)
which show that the energy density for the curvature term does not follow a conservation equation. Moreover the above equation reveals that only when $R$ is a constant one may expect that the effective curvature energy density and pressure will satisfy the hydrodynamic equation of continuity. $R$ can turn out to be a constant when $k=0$ and $H$ is a constant, specifying a de Sitter universe. Using the above result we can write Eq. (34) as:
$\frac{dQ}{dt}+3\left(H+\frac{f^{\prime\prime}\dot{R}}{3f^{\prime}}\right)Q=0\,$, (36)
which shows that if $Q=0$ initially, then $Q$ will always be zero. This shows that if we choose the initial conditions to be satisfying the new Friedmann equation then, while working with the new pressure equation and the continuity equation, the new Friedmann equation will always be satisfied.
The above calculation shows that in the present theory, which is very different from GR, we can ask the same questions we asked before. The questions are related to the redundancy of dynamical equations. The redundancy problem ultimately leads to the special status of the modified Friedmann equation. This workout shows that in $f(R)$ cosmology all the initial conditions must have to satisfy the modified Friedmann equation which acts like a constraint equation.
Fortunately, in the present case there exists a particular conformal transformation, which converts the whole problem into a form where the gravitational dynamics is generated by the Einstein equations in GR. Once we convert the problem into a problem of GR, we can then apply our understanding of the initial value problem in GR to unravel the initial value problem in $f(R)$ theories. As because conformal transformations produces GR like dynamics where Einstein equations dictate gravitational dynamics, our old discussions hold in the conformally transformed frame, called the Einstein frame. In this language the system (or the frame) where the dynamics is governed by the original $f(R)$ equations (as above, in particular the Eqs. (24), (25)) is called the Jordan frame. It is seen that our general discussion about the redundancy of cosmological equations and the special role of the Friedmann equation holds true both in the Jordan frame and in the Einstein frame.